# Global well-posedness of 2D Navier-Stokes with Dirichlet boundary fractional noise

### v1

@@ 4. Interior regularity
As announced below subsection 1.1, we prove Theorem 1.3(2). To this end, we first prove the interior regularity of $w_{g}$ and afterwards the one of $v$ by exploiting the decomposition introduced in subsection 3.2.
@@ 4.1. Stokes equations
We start showing a lemma concerning the relation between the mild and the weak formulation of (1.3) defined below.
Definition 4.1.
$w_{g}\in C(0,T;\mathbb{L}^{r})$
Since $g$ is time-independent, the last term in (4.1) can be rewritten as a stochastic integral as
Lemma 4.2.
Let 1.1 be satisfied. There exists a unique weak solution of (1.3) in the sense of 4.1 and it is given by the formula (3.1).
Proof.
We split the proof into two steps.
Step 1: There exists a unique weak solution of (1.3) and it is necessarily given by the mild formula (3.1). Let $\psi\in C^{1}([0,T];\mathsf{D}(A))$. Arguing as in the first step of the proof of [29, Theorem 1.7], see also [44, Lemma 3], one can readily check that $w_{g}$ satisfies
for each $t\in[0,T],\ \mathbf{P}-a.s.$ The stochastic integral in the relation above is well-defined as a real-valued stochastic integral. Recalling that $(W^{\mathcal{H}}_{t})_{t\geq 0}$ is a $U$-cylindrical fractional Brownian motion we observe that $\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}$ is given by the linear operator on $U$
In conclusion, the stochastic integral in (4.1) is well-defined as a real-valued one, since
Step 2: The mild formula (3.1) is a weak solution of (1.3) in the sense of 4.1. We begin by noticing that $w_{g}$ has the required regularity due to 3.1. Let us test our mild formulation (3.1) against functions $\phi\in\mathsf{D}(A^{2})$. It holds, exploiting self-adjointness property of $S$ and $A$
The relation (4.1) is true. Indeed,
Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution to (3.7) as defined in 3.5 given by Theorem 3.7. Let $N_{0}$ be the $\mathbf{P}$ null measure set where at least one between
@@
(3.10), (3.11), (3.12), (4.1) is not satisfied. In the following, we will work pathwise in $\Omega\setminus N_{0}$ even if not specified. Thanks to the weak formulation guaranteed by 4.2 we can easily obtain the interior regularity of the linear stochastic problem (1.3). Indeed, we are exactly in the same position of [1, Corollary 4.4] and the following holds. We omit the proof as it follows verbatim the one of [1, Corollary 4.4].
Corollary 4.3.
Let 1.1 be satisfied. Let $w_{g}$ be the unique weak solution of (1.3) in the sense of 4.1. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
To deal with the interior regularity of (3.4) we perform a Serrin type argument, see [41, 58]. In contrast to [1], as $w_{g}\notin C([0,T];\mathbb{L}^{4})$, we cannot work directly on $v$. However, recalling that the solution $v$ to (1.2) proven in subsection 3.3 satisfies
@@
where, again, $(v_{1},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.7), cf., subsection 3.2. The advantage of having the splitting (4.7) at our disposal is that $v_{i}$ satisfies a linear problem where the forcing terms only depend on $v_{1},\dots,v_{i-1}$. Thus, by 4.3 and an induction argument, we can prove that $v_{i}$ is smooth inside $(0,T)\times\mathcal{O}$. While to prove the corresponding statement for $\overline{v}$, we can exploit that $\overline{v}$ is a Leray solution (i.e., it has finite energy) and therefore the Serrin regularization can be adjusted to our situation.
We begin with analyzing the interior regularity of $v_{i}$ for $i\in\{1,\dots,N-1\}$.
@@
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5. Then for all $i\in\{0,\dots,N-1\},$ $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
As in the first step of the proof of Theorem 3.7 we argue by induction exploiting strongly the linear and triangle structure of (3.7). Before starting we observe that, by [53, Theorem 3.4.8], it follows that
@@
Then, again by maximal $L^{p}$-regularity techniques for the heat equation (see e.g. [53, Theorem 4.4.4] [37, Theorems 10.2.25 and 10.3.4]) and the trace embedding of [53, Theorem 3.4.8],
@@
From the regularity of $\omega_{0}$, by standard elliptic regularity theory (see for example [61, Chapter 4]), it follows that $\phi_{0}v_{0}\in C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4];H^{1,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\cap L^{p}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4;H^{2,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\ \mathbf{P}-a.s$. Therefore, since $\phi_{0}\equiv 1$ on $B(x_{0},\rho+\varepsilon/8)$
@@
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
As described in 4.2, arguing as in the proof of [29, Theorem 7], we can extend the weak formulation satisfied by $\overline{v}$ to time-dependent test functions $\phi\in C^{1}([0,T];H)\cap C([0,T];\mathsf{D}(A))$ obtaining that for each $t\in[0,T]$
@@
Therefore $h\in L^{2}(0,T;H^{-1}(\mathbb{R}^{2}))+L^{1}(0,T;L^{2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$ Then, arguing as in the first step of the proof of [41, Theorem 13.2], the fact that ${\omega}^{*}$ is a distributional solution of (4.13) implies that ${\omega}^{*}\in C([0,T];L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2})).$ Therefore
@@
By 4.3, 4.4 and relation (4.2) it follows that
@@
This concludes the proof of 4.5. ∎
@@
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.7) in the sense of 3.5. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
and $\omega^{*}=\omega\psi\in L^{2}((0,T)\times\mathbb{R}^{2})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then, arguing as in the proof of 4.5, it follows that $\omega^{*}$ is a distributional solution in $(0,T)\times B(x_{0},\rho+\varepsilon)$ of
@@
From the regularity of $\omega,\ \overline{v},\ \widetilde{\omega},\ \widetilde{v},\ \omega_{N-1},v_{N-1}$, then $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{-1}(B(x_{0},\rho+\varepsilon)))\ \mathbf{P}-a.s$. By standard regularity theory for the heat equation, see for example Step 2 in [41, Theorem 13.1], a solution of (4.17) with $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{k-1}(B(x_{0},\rho+\varepsilon))),\ k\in\mathbb{N}$, belongs to $C([{t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2];H^{k}(B(x_{0},{\rho}+\varepsilon/2)))\cap L^{2}({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2;H^{k+1}(B(x_{0},{\rho}+\varepsilon/2)))$. Therefore
@@
Reiterating the argument as in Step 3 in [41, Theorem 13.1] the claim follows. ∎
@@
The claim follows by 4.3, 4.4 and 4.6 and a localization argument. To begin, recall from the proof of Theorem 1.3(1) in subsection 3.3 that there exists a solution (1.2) on the time interval $[0,T+1]$ and it is given by $\widetilde{u}=w_{g}+\sum_{i=0}^{N-1}v_{i}+\overline{v}$ where $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.7) on $[0,T+1]$ for $r<2q_{\mathcal{H}}$, $N$ as in (3.5) and $p\geq 2^{N}\frac{r}{r-2}$. Then, by 4.3, 4.4, 4.6 and a standard covering argument, for all $t_{0}\in(0,T)$, $\mathcal{O}_{0}\subset\mathcal{O}$ such that $\mathrm{dist}({\mathcal{O}_{0}},\partial\mathcal{O})>0$,

### v2

@@ 4. Interior regularity
As announced at the end of Section 1.1, we prove Theorem 1.3(2). To this end, we first prove the interior regularity of $w_{g}$ and afterwards the one of $v$ by exploiting the decomposition introduced in Section 3.2.
@@ 4.1. Stokes equations
Let $(v_{0},v_{1},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution to (3.13) as defined in Definition 3.7 given by Theorem 3.9. Let $N_{0}$ be the $\mathbf{P}$ null measure set where at least one between
@@
(3.16), (3.17), (3.18), (3.4) is not satisfied. In the following, we will work pathwise in $\Omega\setminus N_{0}$ even if not specified. Thanks to the weak formulation guaranteed by Lemma 3.5 we can easily obtain the interior regularity of the linear stochastic problem (1.4). Indeed, we are exactly in the same position of [1, Corollary 4.4] and the following holds. We omit the proof as it follows verbatim the one of [1, Corollary 4.4].
Lemma 4.1.
Let 1.1 be satisfied. Let $w_{g}$ be the unique weak solution of (1.4) in the sense of Definition 3.4. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
To deal with the interior regularity of (3.10) we perform a Serrin type argument, see [43, 60]. In contrast to [1], as $w_{g}\notin C([0,T];\mathbb{L}^{4})$, we cannot work directly on $v$. However, recalling that the solution $v$ to (1.2) proven in Section 3.3 satisfies
@@
where, again, $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.13), cf. Section 3.2. The advantage of having the splitting (4.1) at our disposal is that $v_{i}$ satisfies a linear problem where the forcing terms only depend on $v_{0},\dots,v_{i-1}$. Thus, by Lemma 4.1 and an induction argument, we can prove that $v_{i}$ is smooth inside $(0,T)\times\mathcal{O}$. While to prove the corresponding statement for $\overline{v}$, we can exploit that $\overline{v}$ is a Leray solution (i.e. it has finite energy) and therefore the Serrin regularization can be adjusted to our situation.
We begin with analyzing the interior regularity of $v_{i}$ for $i\in\{0,\dots,N-1\}$.
@@
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{0},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.13) in the sense of Definition 3.7. Then for all $i\in\{0,\dots,N-1\},$ $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
As in the first step of the proof of Theorem 3.9 we argue by induction exploiting strongly the linear and triangle structure of (3.13). Before starting we observe that, by [55, Theorem 3.4.8], it follows that
@@
Then, again by maximal $L^{p}$-regularity techniques for the heat equation (see e.g. [55, Theorem 4.4.4] [39, Theorems 10.2.25 and 10.3.4]) and the trace embedding of [55, Theorem 3.4.8],
@@
From the regularity of $\omega_{0}$, by standard elliptic regularity theory (see for example [64, Chapter 4]), it follows that $\phi_{0}v_{0}\in C([t_{1}-\varepsilon/4,t_{2}+\varepsilon/4];H^{1,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\cap L^{p}(t_{1}-\varepsilon/4,t_{2}+\varepsilon/4;H^{2,q_{0}}(B(x_{0},\rho+\varepsilon/4);\mathbb{R}^{2}))\ \mathbf{P}-a.s$. Therefore, since $\phi_{0}\equiv 1$ on $B(x_{0},\rho+\varepsilon/8)$
@@
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{0},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.13) in the sense of Definition 3.7. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
As described in Lemma 3.5, arguing as in the proof of [31, Theorem 7], we can extend the weak formulation satisfied by $\overline{v}$ to time-dependent test functions $\phi\in C^{1}([0,T];H)\cap C([0,T];\mathsf{D}(A))$ obtaining that for each $t\in[0,T]$
@@
Therefore $h\in L^{2}(0,T;H^{-1}(\mathbb{R}^{2}))+L^{1}(0,T;L^{2}(\mathbb{R}^{2}))\ \mathbf{P}-a.s.$ Then, arguing as in the first step of the proof of [43, Theorem 13.2], the fact that ${\omega}^{*}$ is a distributional solution of (4.7) implies that ${\omega}^{*}\in C([0,T];L^{2}(\mathbb{R}^{2}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{2})).$ Therefore
@@
By Lemma 4.1, Lemma 4.2 and relation (4.2) it follows that
@@
This concludes the proof of Lemma 4.3. ∎
@@
Let 1.1, $r\in(2,4)$ and $p\geq 2^{N}\frac{r}{r-2}$. Let $(v_{0},\dots,v_{N-1},\overline{v})$ be the $(p,r)$-solution of (3.13) in the sense of Definition 3.7. Then, for all $0<t_{1}\leq t_{2}<T,$ $x_{0}\in\mathcal{O}$, $\rho>0$ such that $\mathrm{dist}({B(x_{0},\rho)},\partial\mathcal{O})>0$,
@@
and $\omega^{*}=\omega\psi\in L^{2}((0,T)\times\mathbb{R}^{2})$ supported in $[t_{1}-\varepsilon,t_{2}+\varepsilon]\times B(x_{0},\rho+\varepsilon)$, then, arguing as in the proof of Lemma 4.3, it follows that $\omega^{*}$ is a distributional solution in $(0,T)\times B(x_{0},\rho+\varepsilon)$ of
@@
From the regularity of $\omega,\ \overline{v},\ \widetilde{\omega},\ \widetilde{v},\ \omega_{N-1},v_{N-1}$, then $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{-1}(B(x_{0},\rho+\varepsilon)))\ \mathbf{P}-a.s$. By standard regularity theory for the heat equation, see for example Step 2 in [43, Theorem 13.1], a solution of (4.11) with $\widetilde{h}\in L^{2}(t_{1}-\varepsilon,t_{2}+\varepsilon;H^{k-1}(B(x_{0},\rho+\varepsilon))),\ k\in\mathbb{N}$, belongs to $C([{t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2];H^{k}(B(x_{0},{\rho}+\varepsilon/2)))\cap L^{2}({t}_{1}-\varepsilon/2,{t}_{2}+\varepsilon/2;H^{k+1}(B(x_{0},{\rho}+\varepsilon/2)))$. Therefore
@@
Reiterating the argument as in Step 3 in [43, Theorem 13.1] the claim follows. ∎
@@
The claim follows by Lemma 4.1, Lemma 4.2 and Corollary 4.4 and a localization argument. To begin, recall from the proof of Theorem 1.3(1) in Section 3.3 that there exists a solution (1.2) on the time interval $[0,T+1]$ and it is given by $\widetilde{u}=w_{g}+\sum_{i=0}^{N-1}v_{i}+\overline{v}$ where $(v_{0},\dots,v_{N-1},\overline{v})$ is the $(p,r)$-solution to (3.13) on $[0,T+1]$ for $r<2q_{\mathcal{H}}$, $N$ as in (3.11) and $p\geq 2^{N}\frac{r}{r-2}$. Then, by Lemma 4.1, Lemma 4.2, Corollary 4.4 and a standard covering argument, for all $t_{0}\in(0,T)$, $\mathcal{O}_{0}\subset\mathcal{O}$ such that $\mathrm{dist}({\mathcal{O}_{0}},\partial\mathcal{O})>0$,